A tree casts a shadow 21 m long. At the same time, the shadow cast by a 46-cm tall statue is 69 cm long. Find the height of the tree. Round results to the nearest unit
Answer by checkley77(12844)

X/21=46/69
69X=21*46
69X=966
X=966/69
X=14 METERS IS THE HEIGHT OF THE TREE WITH THE 21 m. SHADOW.